OpenStudy (narissa):

what is the slope of the line that passes through the points (1,3) and (5,-2) @mazmaz

OpenStudy (narissa):

is the answer 4/5?

OpenStudy (narissa):

@studygurl14

OpenStudy (narissa):

@damirax

OpenStudy (studygurl14):

No, it's not.

OpenStudy (studygurl14):

Slope formula: \(\Large\frac{y_2-y_1}{x_2-x_1}=\frac{-2-3}{5-1}\)

OpenStudy (narissa):

sorry my computer froze is it -4/5?

OpenStudy (narissa):

@newtonson

OpenStudy (narissa):

@random231

OpenStudy (anonymous):

hi

OpenStudy (narissa):

hello is the answer to this question -4/5?

random231 (random231):

nope

OpenStudy (narissa):

-5/4

OpenStudy (anonymous):

no, this -5/4

random231 (random231):

studygurl already answered you narissa

OpenStudy (narissa):

okay thanks :))